Find the exact values of the trigonometric functions of a/2 and 2a by using half-angle and double angle formulas.



1) sin(a) = 4/5, and a in quadrant 1


1
Expert's answer
2021-11-01T18:44:53-0400

1) Use half-angle identity for sine:


"sin \frac{a}{2}=±\sqrt \frac{1-cos a}{2}"

Find "cos a:"


"cos a=±\sqrt{1-sin^2 a}""cos a=±\sqrt{1-(\frac{4}{5})^2}=±\sqrt{1-\frac{16}{25}}=±\sqrt{\frac{9}{25}}=±\frac{3}{5}"

Since "a" is in the quadrant I, "cos a>0, cos a=\frac{3}{5}"

Hence,


"sin \frac{a}{2}=±\sqrt{\frac{1-\frac{3}{5}}{2}}=±\sqrt{\frac{\frac{2}{5}}{2}}=±\sqrt{\frac{1}{5}}"

Since "a" is in the quadrant I, "sin \frac{a}{2}>0, sin \frac{a}{2}=\sqrt{\frac{1}{5}}"

2) Use half-angle identity for cosine:


"cos \frac{a}{2}=±\sqrt \frac{1+cos a}{2}""cos \frac{a}{2}=±\sqrt{\frac{1+\frac{3}{5}}{2}}=±\sqrt{\frac{\frac{8}{5}}{2}}=±\sqrt{\frac{4}{5}}"


Since "a" is in the quadrant I, "cos \frac{a}{2}>0, cos \frac{a}{2}=\sqrt{\frac{4}{5}}"

3) Half-angle for tangent and cotangent:


"tan \frac{a}{2}=sin \frac{a}{2}:cos \frac{a}{2}=\sqrt{\frac{1}{5}}:\sqrt{\frac{4}{5}}=\sqrt{\frac{1}{4}}=\frac{1}{2}""cot \frac{a}{2}=cos \frac{a}{2}:sin \frac{a}{2}=\sqrt{\frac{4}{5}}:\sqrt{\frac{1}{5}}=\sqrt{\frac{4}{1}}=2"

4) Use double-angle identity for sine:


"sin 2a=2sinacos a""sin 2a=2\cdot \frac{4}{5} \cdot \frac{3}{5}=\frac{24}{25}"

5) Use double-angle identity for cosine:


"cos 2a=1-2sin^2a=1-2\cdot({\frac{4}{5}})^2=1-\frac{32}{25}=-\frac{7}{25}"

6) Double-angle for tangent and cotangent:


"tan2a=sin2a:cos 2a=\frac{24}{25}:(-\frac{7}{25})=-\frac{24}{7}"

"cot2a=cos2a:sin 2a=-\frac{7}{25}:\frac{24}{25}=-\frac{7}{24}"

Answer: "sin \frac{a}{2}=\sqrt{\frac{1}{5}}, cos \frac{a}{2}=\sqrt{\frac{4}{5}}, tan \frac{a}{2}=\frac{1}{2}, cot\frac{a}{2}=2,"

"sin 2a=\frac{24}{25}, cos 2a=-\frac{7}{25}, tan 2a=-\frac{24}{7}, cot 2a=-\frac{7}{24}"